# Self-supervised pretraining of vision transformers for animal behavioral analysis and neural encoding

### v1

@@ 4 Results
We demonstrate the versatility of Beast through comprehensive evaluation across three downstream neuro-behavioral tasks: (1) neural encoding, which challenges the model to extract spatiotemporal features that can predict patterns in neural activity; (2) pose estimation, which assesses the model’s ability to extract fine-grained appearance details; and (3) action segmentation, which evaluates the model’s capacity to extract spatiotemporal information required for understanding behavioral sequences. Throughout these evaluations, we present systematic ablation experiments that demonstrate the critical importance of our combined loss functions.
Figure 2: Beast improves neural encoding. A: Example video frame from each dataset. B: Encoding performance is evaluated across multiple baseline features with both linear models (hatched bars; reduced rank regression, or RRR [51]) and nonlinear models (solid bars; temporal convolution network, or TCN [63]). CEBRA [55] uses a contrastive (but not masked autoencoding) loss to embed video frames in a latent feature space. “Motion energy” for the IBL-whisker dataset is a 1D estimate of movement calculated as the sum of the absolute pixel differences between successive frames. Beast features outperform all baselines in both linear and nonlinear regimes. Boxplot showing variability across five test sessions. C: Scatterplot comparison of Beast vs keypoint-based model performance in an example session. Each dot corresponds to an individual neuron. The values in the bottom-right corner represent the session-averaged BPS. D: Top, middle: comparison of the predicted trial-averaged firing rates for Beast and keypoints (lines) and single-trial variability obtained by subtracting the neuron’s average firing rate on each trial (heatmaps). Bottom: comparison of predicted neural principal components for the Facemap dataset.
@@ 4.1 Neural encoding
Predicting neural activity from behavior videos represents a significant challenge with promising implications for understanding the relationship between brain and behavior [44, 58, 71]. Traditional approaches often rely on keypoints [28, 63], potentially missing critical behavioral features that are not included in tracking or are obscured by fur or feathers. While several studies have employed Principal Component Analysis (PCA) [44, 58], this linear technique may inadequately capture subtle behavioral nuances. Vision transformer embeddings offer a compelling alternative, potentially outperforming linear approaches without being constrained by predefined keypoints, thereby providing richer behavioral representations that could reveal previously undetectable neuro-behavioral correlations.
Datasets
@@
We present results on three high-quality neuro-behavioral datasets employing diverse neural recording technologies (Fig. 2). The first dataset is a head-fixed mouse performing a decision-making task from the International Brain Laboratory (IBL) Reporducible Electrophysiology study [28]. This dataset, “IBL,” features simultaneous behavioral video and neural activity monitoring at single-cell, single-spike resolution using Neuropixels probes [21] spanning multiple brain regions (average of 168 neurons per session). The second dataset, “IBL-whisker,” reuses the same sessions as “IBL” but utilizes a cropped area around the whisker pad in the video, a particularly salient behavioral feature for predicting neural activity [58, 74, 63]. The third dataset comes from the Facemap study [63], where neural activity is captured through two-photon calcium imaging, a technique capable of resolving a large number of individual cells, but unable to detect individual spikes. Following the authors’ approach, we predict the principal components of the neural activity to capture predominant variance patterns aross the large recorded neural populations.
Models
@@
We first describe various feature representations used for neural encoding, followed by two models (linear and nonlinear) that we fit to each representation. The first representation for IBL and Facemap are keypoints tracked across the face and body (11 for IBL, 12 for Facemap). For the IBL-whisker dataset, which lacks keypoints, we instead utilize a 1D estimate of whisker pad motion energy (Appendix C). The second representation utilizes PCA applied to raw video frames, while the third leverages CEBRA [55] (which employs a contrastive loss to embed inputs in a latent feature space) applied to raw video frames. Finally, we present results (using the CLS token) from Beast, which is initialized with ImageNet weights then fine-tuned separately on each test session. We train two encoders: linear encoders, which reveal how directly accessible information is within the features, and nonlinear encoders, which better determine the upper bounds of information content in the features. The linear encoder is a reduced rank regression model trained across multiple sessions [81]. The nonlinear encoder is the temporal convolution network proposed in the Facemap study [63].
Evaluation
@@
All model hyperparameters are tuned to ensure robust baseline performance (Appendix C). To evaluate our neural encoding approaches, we utilize the Bits Per Spike (BPS) metric [47] on the spike-resolved IBL dataset (higher values indicate better performance), and the $R^{2}$ metric on the neural principal components in the Facemap dataset. All models are evaluated on five test sessions.
Results
@@
Table 1: Zero-shot neural encoding performance (BPS±$\pm$1 SD).
@@
ViT-M (IN) $0.325\pm 0.091$ $0.307\pm 0.068$
ViT-M (IN+PT) $0.334\pm 0.098$ $0.316\pm 0.073$
Beast (IN+PT) $0.337\pm 0.103$ $0.317\pm 0.083$
Beast (IN+PT+FT) $\mathbf{0.352\pm 0.106}$ $\mathbf{0.335\pm 0.079}$
The ViT-based models in Fig. 2 are fine-tuned individually for each session to enable direct comparison with baseline approaches. We investigated whether pretraining provides additional benefits (Table 1). Strikingly, models pretrained on ImageNet using only the MAE loss, “ViT-M (IN)”, outperform baselines without any fine-tuning. Further pretraining on 77 IBL sessions, “ViT-M (IN+PT)”, improves performance on both IBL datasets, validating the importance of domain-specific pretraining. By incorporating the contrastive objective, Beast achieves superior zero-shot performance: Beast (IN+PT) outperforms the MAE-only variant. Session-specific fine-tuning, “Beast (IN+PT+FT)”, provides additional significant gains, reaching performance levels comparable to models fine-tuned directly from ImageNet weights (Table 11). Notably, even without fine-tuning, domain-specific pretrained models remain highly competitive, offering researchers a practical option when computational resources for fine-tuning are limited. Finally, we experimented with using the patch embeddings as input to the neural encoder, but found superior performance with the CLS tokens (Table 10).
@@ 4.2 Pose estimation
Pose estimation is a fundamental technique in animal behavior analysis [49], enabling precise quantification of posture and movement. Unlike human pose estimation, which benefits from extensive labeled datasets and standardized anatomy, animal pose estimation presents unique challenges such as scarcity of large annotated datasets and significant morphological diversity across species. Pretraining vision transformers on large volumes of unlabeled behavior videos can potentially reduce the labeled data requirements for accurate keypoint localization in various experimental paradigms.
Datasets
@@
We present results on four distinct datasets (Fig. 3): (1) a head-fixed mouse performing a decision-making task from the IBL [29]; (2) a head-fixed mouse running on a treadmill, seen from two views [72]; (3) the Caltech Resident-Intruder Mouse (CRIM13) dataset, consisting of two freely interacting mice [6]; and (4) a freely moving weakly electric fish, seen from three views [3].
@@
Figure 3: Beast improves pose estimation. A: Example frame from each dataset overlaid with ground truth annotations. Green stars indicate the highlighted keypoint in panel B. B: Example traces from the ResNet-50 (gray) and Beast (green) models for a single keypoint in a held-out video. Beast traces evolve more smoothly in time and do not contain erroneous jumps like the ResNet-50 baseline. C: Pixel error as a function of keypoint difficulty (see main text; smaller is better): left-hand side shows performance across all keypoints; moving to the right drops the easier keypoints defined by inter-seed and -model prediction variance. Vertical dashed lines indicate the percentage of data used for the pixel error computation. ViT-M (IN) is a ViT backbone pretrained on ImageNet with a masked autoencoding loss; ViT-M (IN+PT) uses the same architecture and loss but is initialized with ImageNet-pretrained weights then further pretrained on experiment-specific unlabeled frames.
Models
@@
We implemented our pose estimation models using the Lightning Pose framework [3]. We established a strong baseline utilizing a ResNet-50 backbone pretrained on the AP-10K dataset [79], which previous research has shown outperforms ImageNet-pretrained alternatives for animal pose estimation tasks [3]. To evaluate architectural advantages, we incorporated a Vision Transformer (ViT-B/16) pretrained on ImageNet as our second baseline [15], enabling assessment of potential improvements when transitioning from convolutional- to transformer-based architectures. Our own ViT-based models utilize this same architecture. For consistency across all model variants, we employ an identical pose estimation head that transforms backbone features into keypoint heatmaps.
Evaluation
@@
To rigorously evaluate our pose estimation models, we designed a challenging limited-data scenario with only 100 labeled training frames, a realistic constraint for many specialized research settings where extensive annotation is impractical. We measured pixel error between predicted keypoints and ground truth on a test set of novel subjects. For each backbone, we fit three models on different 100-frame subsets. Results are presented as pixel error relative to ensemble standard deviation (e.s.d.) across all seeds and backbones following [3], with error curves showing performance at varying difficulty thresholds. Each point corresponds to keypoints with e.s.d. exceeding the threshold value, with the leftmost portion showing error across all keypoints and rightward movement including only increasingly challenging keypoints (those with higher inter-model variability).
Results
@@
We find robust improvements in pose estimation quality across all datasets when utilizing Beast (Fig. 3). The ImageNet-pretrained ViT outperforms the AP-10K-pretrained ResNet-50 on all datasets except the fish, demonstrating the effectiveness of transformers even with limited labels. Notably, pretraining the transformer with the MAE objective on experiment-specific data yields substantial performance gains across all datasets, including the challenging fish dataset. Augmenting MAE with the contrastive objective (Beast) produces additional performance improvements for some datasets, with particularly pronounced benefits observed in the IBL dataset. We also find these performance advantages persist when scaling to larger training datasets (Fig. 7).
@@
Figure 4: Beast improves action segmentation. A: Example frame from each dataset; action segmentation performance evaluated across multiple baseline features with both TCN (solid) and ensembled (hatched) models. Beast single-model patch embeddings outperform all baselines in the resident-intruder assay, and perform about as well as keypoints in IBL. With ensembling, Beast matches or exceeds all baselines. Error bars on single models represent standard error of the mean across five random initializations. B: Confusion matrices for the single TCN models based on keypoints and Beast patch embeddings. C: Example behavior sequences with feature traces (single seed shown for Beast models), ensemble probabilities, ensembled model-predicted ethograms, ground truth ethograms, and error frames. PCs of SimBA and Beast features are shown for illustration, but the models utilize the full feature set.
@@ 4.3 Action segmentation
Action segmentation classifies discrete behaviors using spatiotemporal video features [49]. Similar to pose estimation, a central challenge in animal action segmentation is the lack of large annotated datasets, as behaviors of interest often vary across species and experimental contexts. Many current approaches rely on keypoints [5, 22, 57, 13, 14], requiring an initial labor-intensive and error-prone preprocessing step. Vision transformer embeddings eliminate this preprocessing requirement and provide an attractive alternative if they match or exceed the performance of keypoint-based methods.
Datasets
@@
We present results on two datasets (Fig. 4): (1) the “IBL” dataset [29], which contains four behavior classes for the paw closest to the camera: still, fidget, wheel-turn, and groom; and (2) the Caltech Mouse Social Interactions (CalMS21) dataset [59], a resident–intruder assay of interacting mouse pairs that contains four social behavior classes: attack, investigation, mount, and other.
Models
@@
We implemented action segmentation models using two types of embeddings from frozen ViT models: (1) CLS token embeddings and (2) per-patch embeddings. For CLS embeddings, we tested both a linear model and a temporal convolution network (TCN) [32], enriching the input with inter-frame differences [4]. We used a sliding window over this feature sequence to predict the action class of the central frame. For patch embeddings, we applied multi-head attention pooling [33, 80, 62] to integrate information across patches, then concatenated the resulting frame-level embeddings with their inter-frame differences before processing them through a TCN (Fig. 8).
@@
For the IBL dataset, we compared against three baseline features: (1) a single paw keypoint, obtained using five pose estimation networks (each trained with 7,000 labeled frames) post-processed with an Ensemble Kalman Smoother [3]; (2) principal components of the video frames; and (3) the CLS and patch embeddings extracted from a frozen-weight ImageNet-pretrained ViT. For CalMS21, we compared against four baseline features: (1) Trajectory Embedding for Behavior Analysis (TREBA) [60], a self-supervised feature extraction method for keypoint trajectories (seven keypoints on each mouse, Fig. 4A); (2) Simple Behavioral Analysis (SimBA) [14], which manually extracts hundreds of features based on spatiotemporal relationships in trajectories of those same keypoints; (3) principal components of video frames; and (4) the CLS and patch embeddings from a frozen-weight ImageNet-pretrained ViT. The pose estimator used for TREBA and SimBA was trained with 15,000 labeled frames [59]. For all baselines except SimBA, we also concatenated inter-frame differences.
Evaluation
@@
All model hyperparameters are tuned to ensure robust baseline performance (Appendix E). We evaluate model performance on a held-out set of animals using the macro-averaged F1 score. For the CalMS21 dataset, following [59], we focused our evaluation on the attack, investigation, and mount classes, averaging the F1 score over these three behaviors. For all feature types and models we train five networks using different random weight initializations. We also report the results of the TCN ensemble by averaging unnormalized logits across seeds before applying softmax.
Results

### v2

@@ 4 Results
We demonstrate the versatility of Beast through comprehensive evaluation across three downstream neuro-behavioral tasks: (1) neural encoding, which challenges the model to extract spatiotemporal features that can predict patterns in neural activity; (2) pose estimation, which assesses the model’s ability to extract fine-grained appearance details; and (3) action segmentation, which evaluates the model’s capacity to extract spatiotemporal features required for predicting behavioral sequences. Throughout these evaluations, we present systematic ablation experiments that demonstrate the critical importance of our combined loss functions, and explore various adaptation strategies, including the use of CLS tokens or patch embeddings from a frozen backbone, as well as end-to-end fine-tuning.
Figure 2: Beast improves neural encoding. A: Example video frame from each dataset. B: Encoding performance is evaluated across multiple baseline features with both linear models (hatched bars; reduced rank regression, RRR) and nonlinear models (solid bars; temporal convolution network, TCN). CEBRA uses a contrastive loss to embed video frames in a latent feature space. “Motion energy” for the IBL-whisker dataset is a 1D estimate of movement calculated as the sum of the absolute pixel differences between successive frames. Beast features outperform all baselines in both linear and nonlinear regimes. Error bars show standard error of the mean (S.E.M.) of Bits per Spike (BPS) across NN=842 neurons from five test sessions (IBL and IBL-whisker) or S.E.M. of variance explained of the principal components of the neural activity across five test sessions (Facemap; see text). C: Scatterplot comparison of Beast vs keypoint-based model performance in an example session. Each dot corresponds to an individual neuron. The values in the bottom-right corner represent the session-averaged BPS. D: Top, middle: comparison of the predicted trial-averaged firing rates for Beast and keypoints (lines) and single-trial variability obtained by subtracting the neuron’s average firing rate on each trial (heatmaps). Bottom: comparison of predicted neural principal components for the Facemap dataset.
@@ 4.1 Neural encoding
Predicting neural activity from behavior videos represents a significant challenge with promising implications for understanding the relationship between brain and behavior [Musall et al., 2019, Stringer et al., 2019, Wang et al., 2025b]. Traditional approaches often rely on keypoints [IBL et al., 2025b, Syeda et al., 2024], potentially missing critical behavioral features that are not included in tracking or are obscured by fur or feathers. Studies have employed Principal Component Analysis on frames [Musall et al., 2019, Stringer et al., 2019], but this linear technique may inadequately capture subtle behavioral nuances. Transformer embeddings offer a compelling alternative, potentially outperforming linear approaches without being constrained by predefined keypoints, thereby providing richer representations that could reveal previously undetectable neuro-behavioral correlations.
Datasets.
@@
We present results on three high-quality neuro-behavioral datasets employing diverse neural recording technologies (Fig. 2). The first dataset is a head-fixed mouse performing a decision-making task from the International Brain Laboratory [IBL et al., 2025a]. This dataset, “IBL,” features simultaneous behavioral video and neural activity monitoring at single-cell, single-spike resolution using Neuropixels probes [Jun et al., 2017] spanning multiple brain regions (average of 168 neurons per session). The second dataset, “IBL-whisker,” reuses the same sessions but utilizes a cropped area around the whisker pad in the video, a particularly salient behavioral feature for predicting neural activity [Stringer et al., 2019, Whiteway et al., 2021, Syeda et al., 2024]. The third dataset comes from the Facemap study [Syeda et al., 2024], where neural activity is captured through two-photon calcium imaging, a technique capable of resolving a large number of individual cells, but unable to detect individual spikes. Following the authors’ approach, we predict the principal components of the neural activity to capture predominant variance patterns across the large recorded neural populations.
Models.
@@
We first describe various feature representations used for neural encoding, followed by two models (linear and nonlinear) that we fit to each representation. The first representation for IBL and Facemap are keypoints tracked across the face and body (11 for IBL, 12 for Facemap). For the IBL-whisker dataset, which lacks keypoints, we instead utilize a 1D estimate of whisker pad motion energy (Appendix C). The second representation utilizes PCA applied to raw video frames, while the third leverages CEBRA [Schneider et al., 2023] (which employs a contrastive loss to embed inputs in a latent feature space) applied to raw video frames. Finally, we present results (using the CLS token) from Beast, which is initialized with ImageNet weights then fine-tuned separately on each test session. Table 13 shows additional baselines that use frozen features from pretrained MAE [He et al., 2022], DINOv2 [Oquab et al., 2023], and CLIP [Radford et al., 2021] models. We train two encoders: linear encoders, which reveal how directly accessible information is within the features, and nonlinear encoders, which better determine the upper bounds of information content in the features. The linear encoder is a reduced rank regression model [Zhang et al., 2025a]. The nonlinear encoder is the temporal convolution network (TCN) proposed in the Facemap study [Syeda et al., 2024].
Evaluation.
@@
All model hyperparameters are tuned to ensure robust baseline performance (Appendix C). To evaluate our neural encoding approaches, we utilize the Bits Per Spike (BPS) metric [Pei et al., 2021] on the spike-resolved IBL dataset (higher values better), and the $R^{2}$ metric on the neural principal components in the Facemap dataset. All models are evaluated on five test sessions.
Results.
@@
[…]
@@
ViT-M (IN) $0.321\pm 0.013$ $0.301\pm 0.012$
ViT-M (IN+PT) $0.331\pm 0.013$ $0.311\pm 0.013$
ViT-C (IN+PT) $0.314\pm 0.013$ $0.283\pm 0.011$
Beast (IN+PT) $0.292\pm 0.012$ $0.309\pm 0.013$
Beast (IN+PT+FT) $\mathbf{0.347\pm 0.014}$ $\mathbf{0.326\pm 0.013}$
The ViT-based models in Fig. 2 are fine-tuned individually for each session to enable direct comparison with baseline approaches. We investigated whether pretraining provides additional benefits (Table 1). Strikingly, models pretrained on ImageNet using only the MAE loss, “ViT-M (IN)”, outperform previous baselines without any fine-tuning. Further pretraining on 77 IBL sessions, “ViT-M (IN+PT)”, improves performance on both IBL datasets, validating the importance of domain-specific pretraining. By incorporating the contrastive objective, Beast achieves superior zero-shot performance: Beast (IN+PT) outperforms the MAE-only variant, as well as a contrastive-only variant ViT-C. Session-specific fine-tuning, “Beast (IN+PT+FT)”, provides additional significant gains, reaching performance levels comparable to models fine-tuned directly from ImageNet weights (Table 13). Notably, even without fine-tuning, domain-specific pretrained models remain highly competitive, offering researchers a practical option when computational resources for fine-tuning are limited. Finally, we experimented with using the patch embeddings as input to the neural encoder, but found superior performance with the CLS tokens (Table 12).
@@ 4.2 Pose estimation
Pose estimation is a fundamental technique in animal behavior analysis [Pereira et al., 2020], enabling precise quantification of posture and movement. Unlike human pose estimation, which benefits from extensive labeled datasets and standardized anatomy, animal pose estimation presents unique challenges such as scarcity of large annotated datasets and significant morphological diversity across species. Pretraining models on large volumes of unlabeled behavior videos can potentially reduce the labeled data requirements for accurate keypoint localization in various experimental paradigms.
Datasets.
@@
We present results on four distinct datasets (Fig. 3): (1) a head-fixed mouse performing a decision-making task [IBL et al., 2025a]; (2) a head-fixed mouse running on a treadmill, seen from two views [Warren et al., 2021]; (3) the Caltech Resident-Intruder Mouse (CRIM13) dataset, consisting of two freely interacting mice [Burgos-Artizzu et al., 2012]; and (4) a freely moving weakly electric fish, seen from three views [Biderman et al., 2024, Pedraja et al., 2025].
@@
Figure 3: Beast improves pose estimation. A: Example frame from each dataset overlaid with ground truth annotations. Green stars indicate the highlighted keypoint in panel B. B: Example traces from the ResNet-50 (gray) and Beast (green) models for a single keypoint in a held-out video. Beast traces evolve more smoothly in time and do not contain erroneous jumps like the ResNet-50 baseline. C: Pixel error as a function of keypoint difficulty (see main text; smaller is better): left-hand side shows performance across all keypoints; moving to the right drops the easier keypoints defined by inter-seed and -model prediction variance. Vertical dashed lines indicate the percentage of data used for the pixel error computation. ViT-M (IN) is a ViT backbone pretrained on ImageNet with a masked autoencoding loss; ViT-M (IN+PT) uses the same architecture and loss but is initialized with ImageNet-pretrained weights then further pretrained on experiment-specific unlabeled frames. ViT-C (IN+PT) performs the experiment-specific pretraining using the temporal contrastive loss only.
Models.
@@
We implemented pose estimation models using Lightning Pose [Biderman et al., 2024]. We established a strong baseline utilizing a ResNet-50 backbone pretrained on AP-10K [Yu et al., 2021], which outperforms a DeepLabCut baseline (ImageNet-pretrained ResNet-50) on all but the CRIM13 dataset (Fig. 11). Our second baseline is a Vision Transformer (ViT-B/16) pretrained on ImageNet [He et al., 2022] using our own implementation of ViTPose [Xu et al., 2022], enabling assessment of potential improvements when transitioning from convolutional- to transformer-based architectures. Our own ViT-based models utilize this same architecture. In the Appendix we provide additional baselines that use fine-tuned DINO [Caron et al., 2021], DINOv2 [Oquab et al., 2023], and Segment Anything [Kirillov et al., 2023] encoders (which Beast consistently outperforms; Fig. 9). For consistency across all model variants, we employ an identical pose estimation head that transforms backbone features into keypoint heatmaps. Given the spatial nature of the task, we use patch embeddings rather than CLS tokens in the transformers, and train all models end-to-end.
Evaluation.
@@
To rigorously evaluate our pose estimation models, we designed a challenging limited-data scenario with only 100 labeled training frames, a realistic constraint for many research settings where extensive annotation is impractical. We measured pixel error between predicted keypoints and ground truth on a test set of novel subjects. For each backbone, we fit three models on different 100-frame subsets. Results are presented as pixel error relative to ensemble standard deviation (e.s.d.) across all seeds and backbones following Biderman et al. [2024], with error curves showing performance at varying difficulty thresholds. Each point corresponds to keypoints with e.s.d. exceeding the threshold value, with the leftmost portion showing error across all keypoints and rightward movement including only increasingly challenging keypoints (those with higher inter-model variability).
Results.
@@
We find robust improvements in pose estimation quality across all datasets when utilizing Beast (Fig. 3). The ImageNet-pretrained ViT outperforms the AP-10K-pretrained ResNet-50 on all datasets except the fish, demonstrating the effectiveness of transformers even with limited labels. Notably, pretraining the transformer with the MAE objective on experiment-specific data yields substantial performance gains across all datasets, including the challenging fish dataset. Augmenting MAE with the contrastive objective (Beast) produces additional performance improvements for some datasets, with particularly pronounced benefits observed in the IBL dataset. However, pretraining only on the contrastive objective leads to significantly worse results, consistent with the different learning objectives: the temporal contrastive loss emphasizes high-level temporal structure, whereas the MAE loss emphasizes low-level, pixel-level features. Consequently, MAE-pretrained representations are better suited for pixel-level prediction tasks like pose estimation, though addition of the contrastive loss in Beast still provides complementary benefits. We also find Beast’s performance advantages persist when scaling to larger training datasets (Fig. 10).
@@
Figure 4: Beast improves action segmentation. A: Example frame from each dataset; performance evaluated across multiple baseline features with both TCN (solid) and ensembled (hatched) models. Error bars represent standard error of the mean across five random initializations. B: Confusion matrices for TCN models based on keypoints and Beast patch embeddings. C: Example behavior sequences with feature traces (single seed shown for Beast models), ensemble probabilities, ensembled model ethograms, ground truth ethograms, and error frames. PCs of SimBA and Beast features are shown for illustration, but the models utilize the full feature set.
@@ 4.3 Action segmentation
Action segmentation classifies discrete behaviors using spatiotemporal video features [Pereira et al., 2020]. Similar to pose estimation, a central challenge in animal action segmentation is the lack of large annotated datasets, as behaviors of interest often vary across species and experimental contexts. Many current approaches rely on keypoints [Branson et al., 2009, Kabra et al., 2013, Segalin et al., 2021, Gabriel et al., 2022, Goodwin et al., 2024], requiring an initial labor-intensive and error-prone preprocessing step. Vision transformer embeddings eliminate this preprocessing requirement and provide an attractive alternative if they match or exceed the performance of keypoint-based methods.
Datasets.
@@
We present results on two datasets (Fig. 4): (1) the “IBL” dataset [IBL et al., 2025a], which contains four behavior classes for the paw nearest the camera; and (2) the Caltech Mouse Social Interactions (CalMS21) dataset [Sun et al., 2021a], a resident–intruder assay of interacting mouse pairs that contains four social behavior classes.
Models.
@@
We implemented models using two types of embeddings from frozen ViT models: (1) CLS tokens and (2) per-patch embeddings. For CLS embeddings, we tested both a linear model and a TCN [Lea et al., 2016], enriching the input with inter-frame differences [Blau et al., 2024]. We used a sliding window over this feature sequence to predict the action class of the central frame. For patch embeddings, we applied multi-head attention pooling [Lee et al., 2019, Yu et al., 2022, Sun et al., 2024] to integrate information across patches, then concatenated the resulting frame-level embeddings with their inter-frame differences before processing them through a TCN (Fig. 12).
@@
For IBL, we compared against three baseline features: (1) a single paw keypoint, obtained using five pose estimation networks (each trained with 7,000 labeled frames) post-processed with an Ensemble Kalman Smoother [Biderman et al., 2024]; (2) principal components of the video frames; and (3) the CLS and patch embeddings extracted from a frozen-weight ViT pretrained on ImageNet with MAE loss (which outperformed patch embeddings from DINOv2 in both datasets; Table 15). For CalMS21, we compared against four baseline features: (1) Trajectory Embedding for Behavior Analysis (TREBA) [Sun et al., 2021b], a self-supervised feature extraction method for keypoint trajectories; (2) Simple Behavioral Analysis (SimBA) [Goodwin et al., 2024], which extracts hundreds of hand-crafted features from the keypoint trajectories; (3) principal components of video frames; and (4) the CLS and patch embeddings from a frozen-weight ViT pretrained on ImageNet with MAE loss. The pose estimator used for TREBA and SimBA was trained with 15,000 labeled frames [Sun et al., 2021a]. For all baselines except SimBA, we also concatenated inter-frame differences.
Evaluation.
@@
All model hyperparameters are tuned to ensure robust baseline performance (Appendix E). We evaluate performance on held-out animals using the macro-averaged F1 score. For the CalMS21 dataset, following Sun et al. [2021a], we average the F1 score over the attack, investigation and mount classes. For all models we train five networks using different random initializations. We also report the results of model ensembles by averaging logits across seeds before applying softmax.
Results.